OpenStudy (anonymous):

Jon's new mini computer monitor has a 20 cm by 24 cm LCD screen with a plastic trim border. The total area of the monitor and border is 672 cm2. How wide is the trim around the monitor?

OpenStudy (anonymous):

\[(20+w)(24+w)=672\] \[480+44w+w^2=672\] \[w^2 + 44w - 192 = 0\] \[(w+48)(w-4) = 0\] \[w = -48\] or \[w = 4\] w > 0 so w = 4 the trim is half the extra part wide, so the trim is 2 cm wide!

OpenStudy (anonymous):

I mean its 4 not 2 :D

OpenStudy (anonymous):

Thaaank youu sooo muchhhh'<3